OpenStudy (anonymous):

Say you have a very long string and want the hex nuts to hit the ground 1 second apart. Using the kinematics equations, determine the spacing for 5 nuts to hit with equal timing. How much string would you need?

OpenStudy (anonymous):

d=.5*9.8*t^2 That is your basic equation so d1= .5*9.8*(t+1)^2 so d2= .5*9.8*(t+2)^2 so d3= .5*9.8*(t+3)^2 so d4= .5*9.8*(t+4)^2 so d5= .5*9.8*(t+5)^2 also this is assuming the string is straight up and down. Some sine will have to be include or adjust for acceleration =9.8